OpenStudy (anonymous):

How do you solve S=1/3(x+b)a. By solving equations for x?

mathslover (mathslover):

S=1/3(x+b)a \[S=(x+b)(a)/3\] \[3S=(x+b)(a)\] \[3/a=x+b\] \[x=(3/a)-b\]

OpenStudy (anonymous):

Where has S gone???

OpenStudy (anonymous):

\[x = (1-3aSb)/3aS\] is the answer I think..

OpenStudy (lgbasallote):

\[S = \frac{1}{3(x+b)a}\] \[3Sa(x+b) = 1\] \[3Sax + 3Sab = 1\] \[3Sax = 1 - 3Sab\] \[x = \frac{1 - 3Sab}{3Sa}\] just viewing the possibilities

OpenStudy (lgbasallote):

\[S = \frac{1}{3(x+b)}a\] \[3S(x+ b) = a\] \[3Sx + 3Sb = a\] \[3Sx = a - 3Sb\] \[x = \frac{a - 3Sb}{3S}\]

OpenStudy (lgbasallote):

another possibility lol